Question Two (40 Marks): School Days Furniture, Inc., manufactures a variety of desks, chairs, tables, and shelf units whichB Production budget (in sets) (15 points): July August September Sales 10,000 Add: Desired ending inventory. 2,400 3,000 TotaC Raw matenal purchases (15 points) July August September Planned production (sets) 10,400 Raw material required per set (boaplease make sure the answer is correct 100%
Answer #1

Requirement A:

Sales Budget
July August September
Sales ( in Sets)              10,000              12,000              15,000
Sales price per set $                 40 $                 40 $                 40
Sales revenue $       400,000 $       480,000 $       600,000

Requirement B:

Production Budget
July August September
Sales              10,000              12,000              15,000
Add: Desired ending inventory                2,400                3,000                3,000
Total requirements              12,400              15,000              18,000
Less: projected beginning inventory                2,000                2,400                3,000
Planned production              10,400              12,600              15,000
Note: Desired ending inventory of august = 15000*20% = 3000

Requirement C:

Raw materials purchases
July August September
Planned production (sets)              10,400              12,600              15,000
Raw material required per set (Board feet)                      10                      10                      10
Raw material required for production(board feet)            104,000            126,000            150,000
Add: Desired ending inventory RM (board feet)              12,600              15,000              16,000
Total requirements            116,600            141,000            166,000
Less: projected beginning inventory of RM (board feet)              10,400              12,600              15,000
Planned purchases of raw materials (board feet)            106,200            128,400            151,000
Cost per board foot $              0.60 $              0.60 $              0.60
Planned purchases of raw materials (dollars) $          63,720 $          77,040 $          90,600
Note: Desired ending inventory RM (board feet) of august = 150000*10% = 15000